POLAR (ACYCLIC) ISOMER OF FORMIC ACID DIMER: RAMAN SPECTROSCOPY STUDY

\underline{ROMAN M. BALABIN}, Department of Chemistry and Applied Biosciences, ETH Zurich, 8093 Zurich, Switzerland.

Formic (methanoic) acid spectral range of 575–1150 cm^{−1} has been studied by gas-phase Raman spectroscopy method. A weak Raman-active vibration of polar (acyclic) HCOOH dimer has been found at 864±2.1 cm^{−1} and assigned using quantum chemistry data. The temperature-dependence of intensity ratios of Raman lines was used to evaluate the thermodynamic parameters of polar dimer. Its experimental dimerization enthalpy (△H) was found to be -8.6±0.2 kcal mol^{−1}. Entropy of dimerization has been evaluated using theoretical (MP2) Raman scattering activities. Its value (△S) is estimated as -36±2 cal mol^{−1} K^{−1}. The results are compared with the published experimental data and calculations. The presented results can be used for molecular dynamics simulations, hydrogen bond energy estimation, and analysis of CH_{2}O_{2} vapor density measurements.